Your new job now requires you to make 1.5 L of a solution that is 0.100M potassium phosphate. You find a bottle on the shelf that is 3M potassium phosphate. Describe how you would make that solution.
Answer #1

Moles = concentration × volume

Moles required to make= 0.1M × 1.5L = 0.15 mole

Concentration we have = 3M

Volume we require of 3M = mole/concentration

Volume = 0.15/3 = 0.05L

Thus we have to take 0.05L of 3M solution and dilute it upto 1.5L. then we will obtain 1.5L of 0.1M solution.